This is a similar polygon question involving a pentagon. I couldn't find the topic in Algebra but it's from an algebra class.
Need to find x of the segment vw
The two pentagons are similar. The first one the the sides for Q to R is 4 and R to S is 3 and S to T is 5 and T to P is 3 and P to Q is 2. These sides will somehow apply with the second pentagon. The sides that are suppose to be the comparison for the VW I think is the PQ side. The other sides for the second Pentagon that they want the first compared to and the side needs to be solved for has points of W to X 6.4 and X to Y 4.8 and Y to Z is 8 and Z to V is 4.8.
Please help me solve the VW. I need to get this answer ASAP tonight. I have been up most of the night trying to figure it out.

Answer by Edwin McCravy(20086)

This is a similar polygon question involving a pentagon. I couldn't find the
topic in Algebra but it's from an algebra class.

Need to find x of the segment vw

The two pentagons are similar. The first one the the sides for Q to R is 4 and
R to S is 3 and S to T is 5 and T to P is 3 and P to Q is 2. These sides will
somehow apply with the second pentagon. The sides that are suppose to be the
comparison for the VW I think is the PQ side. The other sides for the second
Pentagon that they want the first compared to and the side needs to be solved
for has points of W to X 6.4 and X to Y 4.8 and Y to Z is 8 and Z to V is 4.8.
Please help me solve the VW. I need to get this answer ASAP tonight. I have
been up most of the night trying to figure it out.

Since pentagon PQRST is similar to pentagon VWXYZ, then

WX is to QR as XY is to RS, or
WX/QR = XY/RS
6.4/4 = 4.8/3
  1.6 = 1.6 

So if these pentagons are similar, then every side of pentagon VWXYZ is
1.6 times its corresponding side of pentagon PQRST.  Let's check to make
sure the given pentagons are really similar:

YZ should then be 1.6 times ST.  Sure enough, 8 = 1.6(5)

and

ZV should then be 1.6 times TP.  Sure enough, 4.8 = 1.6(3)

So VW must be 1.6 times PQ.  So VW = 1.6(PQ) = 1.6(2) = 3.2

That's the answer.  VW = 3.2

Edwin